# On local Galois deformation rings: generalised reductive groups

### v1

@@ 8.2. Finite maps
By topological Nakayama’s lemma it is enough to show, that the fibre ring $C:=k\otimes_{R^{\mathrm{ps}}_{H}}R^{\mathrm{ps}}_{G}$ is a finite-dimensional $k$-vector space. Since $\Gamma$ is topologically finitely generated, we know by [44, Theorem 5.7], that $R^{\mathrm{ps}}_{G}$ is noetherian, hence $C$ is a complete local noetherian ring and we are left to show, that $C$ has Krull dimension 00.
If $C$ is not 00-dimensional then there is a point $x\in\operatorname{Spec}(C)$ of dimension $1$. The residue field $\kappa(x)$ is a local field of characteristic $p$ by [8, Lemma 3.17] and the corresponding homomorphism $C\to\kappa(x)$ gives a $G$-pseudocharacter $\Theta\in\mathrm{PC}_{G}^{\Gamma}(\kappa(x))$, such that $\varphi\circ\Theta\in\mathrm{PC}_{H}^{\Gamma}(\kappa(x))$ arises by scalar extension from the residual pseudocharacter attached to $\varphi\circ\overline{\rho}$.
For all $m\geq 0$, the map $(\varphi\circ\Theta)_{n}=\Theta_{n}\circ\varphi^{*}$ is the composition
$\mathcal{O}[H^{m}]^{H^{0}}\xrightarrow{\varphi^{*}}\mathcal{O}[G^{m}]^{G^{0}}\xrightarrow{\Theta_{n}}\mathrm{Map}(\Gamma^{m},\kappa(x))$
and has image in $\mathrm{Map}(\Gamma^{m},k)$. By 6.3 $\varphi^{*}$ is finite, so $\Theta_{n}$ has image in $\mathrm{Map}(\Gamma^{m},\overline{k})$. This implies that $\kappa(x)$ is algebraic over $k$ and contradicts the assumption that $x$ is a dimension $1$ point. ∎
@@
[…]
@@ 8.3. Products
Recall that a $G$-valued pseudocharacter is a family of $\mathcal{O}$-algebra homomorphisms $\Theta_{m}:\mathcal{O}[G^{m}]^{G^{0}}\to\mathrm{Map}(\Gamma^{m},A)$ satisfying certain compatibility conditions. Restriction along the map $\mathcal{O}[G_{i}^{m}]^{G_{i}^{0}}\to\mathcal{O}[G^{m}]^{G^{0}}$ defines $G_{i}$-pseudocharacters $\Theta^{(i)}$ satisfying the same compatibility conditions; this can be directly seen from the description of pseudocharacters in terms of functors on the category of finitely generated free groups [44, Proposition 3.19]. Conversely given $\Theta^{(i)}\in\mathrm{PC}_{G_{i}}^{\Gamma}(A)$, we can define a $G$-pseudocharacter $\Theta^{(1)}\otimes\Theta^{(2)}$ by defining $(\Theta^{(1)}\otimes\Theta^{(2)})_{m}:\mathcal{O}[G^{m}]^{G^{0}}\to\mathrm{Map}(\Gamma^{m},A)$ as the tensor product $\Theta^{(1)}_{m}\otimes\Theta^{(2)}_{m}:\mathcal{O}[G_{1}^{m}]^{G_{1}^{0}}\otimes_{\mathcal{O}}\mathcal{O}[G_{2}^{m}]^{G_{2}^{0}}\to\mathrm{Map}(\Gamma^{m},A)$. This implies that the first map is bijective.
@@ 8.4. Functoriality in $\Gamma$
If $G=\operatorname{GL}_{d}$ then this is proved in [8, Lemma A.3]. In the general case, one could either use that $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\rightarrow R^{\mathrm{ps}}_{G}$ is finite, which follows from Propositions 6.4 and 7.4, or observe that the proof of [8, Lemma A.3] carries over to the more general setting. ∎
@@ 9.1. Trimming the component group
Since $G$ is generalised reductive $G/G^{0}$ is finite étale over $\mathcal{O}$ by Remark 2.6. We may think of finite étale $\mathcal{O}$-groups schemes as finite groups together with $\operatorname{Gal}(\overline{k}/k)$-action by group automorphisms. Since $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}(k)$ is non-empty, $\Delta$ is contained in $(G/G^{0})(k)$ and hence the constant group scheme $\underline{\Delta}$ defined by $\Delta$ is a finite étale subgroup scheme of $G/G^{0}$.
@@
(39) $\nu:X^{\mathrm{gen},\tau}_{G}(A)\rightarrow(G/G^{0})_{X^{\mathrm{ps}}_{\operatorname{GL}_{d}}}^{N}(A),\quad\rho\mapsto(\pi_{G}(\rho(\gamma_{1})),\ldots,\pi_{G}(\rho(\gamma_{N}))).$
Since $\nu$ is $G^{0}$-equivariant it factors through $X^{\mathrm{git},\tau}_{G}\rightarrow(G/G^{0})^{N}_{X^{\mathrm{ps}}_{\operatorname{GL}_{d}}}$. We view an $N$-tuple $\underline{\delta}=(\delta_{1},\ldots,\delta_{N})\in\Delta^{N}$ as a closed constant subscheme of $\underline{\Delta}^{N}_{X^{\mathrm{ps}}_{\operatorname{GL}_{d}}}$ over $X^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ and define
@@
[…]
@@ 9.2. Absolutely irreducible locus
We say that a geometric point $x\in X^{\mathrm{gen}}_{G}$ is $G$-stable, if the orbit $G\cdot x$ is closed and $Z_{G}(x)^{0}=Z(G)^{0}_{\kappa(x)}$, where $Z_{G}(x)$ is the $G_{\kappa(x)}$-centraliser of $x$.
@@
The action of $G$ on $X^{\mathrm{gen}}_{G}$ factors through the action of $G/Z(G)$. The stability condition in the definition is equivalent to $G\cdot x$ closed and $\dim G\cdot x=\dim(G/Z(G))_{\kappa(x)}$. Hence, it is equivalent to $G/Z(G)$-stability in the sense of [52, Definition 1].
@@
Moreover, $H$ is reductive and $Z_{G}(H)^{0}=Z(G_{\kappa(x)})^{0}$.
@@
The equivalence of (1) and (2) follows from Corollary 6.11. The equivalence of (2) and (3) follows from Definition 2.26. The equivalence of (3) and (4) follows from [46, Proposition 16.7]. The last part follows from [40, Lemma 6.2]. ∎
@@
where the inequality in the middle follows from [40, Lemma 6.2], as the centre of $H_{\kappa(x)}$ is contained in the $L_{\kappa(x)}$-centraliser of $H_{\kappa(x)}$, and the other two follow from Lemma 2.2. ∎
@@
Since extending scalars from $\kappa(y)$ to its algebraic closure does not change the dimension, we may assume that $y$ is a geometric point of $U_{GG}$ and denote its residue field by $\kappa$.
Let $x$ be any closed point in the fibre $X^{\mathrm{gen}}_{G,y}$. Then the $G$-orbit $G\cdot x$ is closed, as $x$ is $G$-stable by Proposition 9.7. Proposition 6.7 implies that the $G^{0}$-orbit $G^{0}\cdot x$ is also closed. Since $X^{\mathrm{gen}}_{G,y}$ contains a unique closed $G^{0}$-orbit by Proposition 6.9 we conclude that $G^{0}(\kappa)$ acts transitively on $X^{\mathrm{gen}}_{G,y}(\kappa)$. Thus the orbit map
$G^{0}_{\kappa}/Z_{G^{0}}(x)\rightarrow X^{\mathrm{gen}}_{G,y}$
induces a bijection on $\kappa$-points. Since both schemes are of finite type over $\kappa$, we deduce that $\dim X^{\mathrm{gen}}_{G,y}=\dim G^{0}_{\kappa}/Z_{G^{0}}(x)=\dim G^{0}_{\kappa}-\dim Z_{G^{0}}(x)$. By passing to neutral components we obtain $Z_{G^{0}}(x)^{0}=Z_{G}(x)^{0}=Z(G)_{\kappa}^{0}$, where the last equality holds as $x$ is $G$-stable. Thus $\dim Z_{G^{0}}(x)=\dim Z(G)_{\kappa}$ and we conclude that
@@
Corollary 9.9 says that $\dim V_{y}=d$, where $d:=\dim G_{k}-\dim Z(G_{k})$, for all $y\in U$. The inequality $\dim V\leq\dim U+d$ follows directly from [8, Lemma 3.18 (6)] (or alternatively [54, Tag 0BAG]).
To prove the reverse inequality we may assume that $U$ is irreducible and, since $V\rightarrow U$ is surjective by [52, Theorem 3 (ii)], we may replace $V$ by an irreducible component, whose image contains the generic point of $U$. Lemma 2.1 in [8] implies that $V$ is $G^{0}$-invariant and [52, Theorem 3 (iii)] implies that the image of $V$ in $U$ is closed, and hence $V\rightarrow U$ is surjective. It follows from [54, Tag 02JU] applied with $s\in U$ a closed point such that $\dim U=\dim\mathcal{O}_{U,s}$ and $x\in V$ a closed point, which maps to $s$, that $\dim V\geq\dim U+d$. ∎
@@ 10. Bounds for the dimensions of reductive subgroups
Let $G$ be a generalised reductive group over an an algebraically closed field $\kappa$. If $H$ is a closed reduced subgroup scheme of $G$, such that the unipotent radical of $H^{0}$ is trivial then $H$ is also generalised reductive by 2.7. Let $G^{\prime}$ denote the derived subgroup of $G^{0}$. Then $G^{\prime}$ is semisimple and $G^{0}=Z(G^{0})G^{\prime}$, [11, 14.2]. Moreover, $G^{\prime}$ is equal to its own derived subgroup, [17, Example 1.1.16].
@@
After replacing $G$ by $G^{\prime}$ we may assume that $G$ is connected, reductive and is equal to its own derived subgroup. One may check that the unipotent radical of $(G^{\prime}\cap H)^{0}$ is normal in $H^{0}$. Since $H^{0}$ is reductive, $(G^{\prime}\cap H)^{0}$ is also reductive. Thus after replacing $H$ by $H\cap G^{\prime}$ we may assume that $H$ is closed reductive subgroup of $G$. Let us assume that $H\neq G$. Then we have to show that $\dim G/H>1$.
There exists a finite dimensional representation $V$ of $G$ and a $1$-dimensional subspace $\ell$ of $V$ such that $H$ is equal to the $G$-stabiliser of $\ell$ in $\mathbb{P}(V)$, [33, 11.2]. Moreover, the $G$-action on $\mathbb{P}(V)$ induces an isomorphism of homogeneous spaces between $G/H$ and the $G$-orbit of $\ell$ in $\mathbb{P}(V)$, [33, 12.1].
Since $H$ is reductive $G/H$ is affine, [45], and since $\dim G/H\geq 1$, $G\cdot\ell$ cannot be closed in $\mathbb{P}(V)$. Its closure contains a closed $G$-orbit $G\cdot\ell^{\prime}$ of strictly smaller dimension, [33, 8.3].
If $\dim G\cdot\ell^{\prime}\geq 1$ then $\dim G\cdot\ell\geq 2$ and we are done. If $\dim G\cdot\ell^{\prime}=0$ then since $G$ is connected we deduce that $G$ fixes $\ell^{\prime}$. Let $v$ be a basis vector for $\ell$ and $v^{\prime}$ a basis vector of $\ell^{\prime}$. Since $G$ is equal to its derived subgroup we deduce that $G$ fixes $v^{\prime}$. Let $\bar{\ell}$ be the image of $\ell$ in $V/\ell^{\prime}$. Since $G\neq H$, $\ell\neq\ell^{\prime}$ and hence $\bar{\ell}$ is $1$-dimensional. Let $\bar{H}$ be the $G$-stabiliser of $\bar{\ell}$ in $\mathbb{P}(V/\ell^{\prime})$. Then $H\subseteq\bar{H}\subseteq G$.
If $\dim G=\dim\bar{H}$ then $\dim G/\bar{H}=0$ and since $G$ is connected $G=\bar{H}$. Since $G$ is equal to its derived subgroup it fixes $v+\ell^{\prime}$, and thus $gv=v+a(g)v^{\prime}$ for all $g\in G$. Since $G$ acts trivially on $v^{\prime}$, $a$ defines a group homomorphism $a:G\rightarrow\mathbb{G}_{a}$. Since $G$ is equal to its own derived subgroup we get that $a=0$ and hence $G$ fixes $v$, which yields a contradiction as the $G$-stabiliser of $v$ is $H$. Thus $\dim\bar{H}<\dim G$.
If $\dim H<\dim\bar{H}$ then we obtain $\dim G-\dim H\geq 2$ and we are done.
If $\dim H=\dim\bar{H}$ then $H^{0}=\bar{H}^{0}$, hence $\bar{H}$ is reductive. We may now repeat the argument with $V/\ell^{\prime}$ instead of $V$. Since $V$ is finite dimensional this process has to terminate eventually yielding $\dim G-\dim H\geq 2$. ∎
@@
If $\dim T=1$ then $G^{\prime}=\operatorname{SL}_{2}$ by [53, 7.2.4], as the assumption that $G^{\prime}$ is simply connected rules out $\operatorname{PGL}_{2}$. If $\mathrm{char}(\kappa)\neq 2$ then the action of $G^{\prime}$ on $\operatorname{Lie}G^{\prime}$ is an irreducible representation, isomorphic to $\operatorname{Sym}^{2}(\mathrm{Std})$, where $\mathrm{Std}$ is the standard $2$-dimensional representation of $\operatorname{SL}_{2}$. If $\mathrm{char}(\kappa)=2$ then $\operatorname{Lie}G^{\prime}$ is a non-split extension $0\rightarrow\mathbf{1}\rightarrow\operatorname{Lie}G^{\prime}\rightarrow V\rightarrow 0$, with $\mathbf{1}$ equal to the subspace of scalar matrices and $V$ an irreducible $2$-dimensional representation. In both cases we obtain the claim.
Let $\dim T$ be arbitrary and let $\chi:\operatorname{Lie}G^{\prime}\rightarrow\mathbf{1}$ be $G^{\prime}$-invariant. Let us abbreviate $\mathfrak{g}=\operatorname{Lie}G^{\prime}$. We have $\mathfrak{g}=\mathfrak{g}^{T}\oplus\bigoplus\nolimits_{\alpha\in\Phi}\mathfrak{g}_{\alpha}$, where $\Phi$ is the set of roots, and $\mathfrak{g}^{T}=\operatorname{Lie}(T)$, [11, 13.18(1)]. Let $\alpha\in\Phi$ and let $G_{\alpha}$ be the centraliser of $(\operatorname{Ker}\alpha)^{0}$ in $G^{\prime}$. Then $G_{\alpha}$ is reductive of semisimple rank $1$, [11, 13.18(4)]. Let $G^{\prime}_{\alpha}$ be the derived subgroup of $G_{\alpha}$. The assumption that $G^{\prime}$ is simply connected implies that $G^{\prime}_{\alpha}\cong\operatorname{SL}_{2}$ by [17, 5.3.9]. It follows from [17, 1.2.7] that $\operatorname{Lie}G_{\alpha}^{\prime}=\mathfrak{g}_{-\alpha}\oplus\operatorname{Lie}(\alpha^{\vee}(\mathbb{G}_{m}))\oplus\mathfrak{g}_{\alpha}$. The restriction of $\chi$ to $\operatorname{Lie}G^{\prime}_{\alpha}$ is $G^{\prime}_{\alpha}$-invariant and hence zero by the previous part. Hence, $\chi$ is zero on $\operatorname{Lie}(\alpha^{\vee}(\mathbb{G}_{m}))$ and $\mathfrak{g}_{\alpha}$ for every coroot $\alpha^{\vee}$.
@@
Let $P$ be an $R$-parabolic subgroup of $G$ and let $L$ be its Levi subgroup and $U$ its unipotent radical. Then $P^{0}=P\cap G^{0}$ is a parabolic subgroup of $G^{0}$ and $L^{0}=L\cap G^{0}$ is a Levi subgroup of $P^{0}$, [40, Proposition 5.2]. In particular, $L^{0}$ is reductive. We let $L^{\prime}$ be the derived subgroup of $L^{0}`$.

### v2

@@ 8.2. Finite maps
By topological Nakayama’s lemma it is enough to show, that the fibre ring $C:=k\otimes_{R^{\mathrm{ps}}_{H}}R^{\mathrm{ps}}_{G}$ is a finite-dimensional $k$-vector space. Since $\Gamma$ is topologically finitely generated, we know by [50, Theorem 5.7], that $R^{\mathrm{ps}}_{G}$ is noetherian, hence $C$ is a complete local noetherian ring and we are left to show, that $C$ has Krull dimension 0.
If $C$ is not 0-dimensional then there is a point $x\in\operatorname{Spec}(C)$ of dimension $1$. The residue field $\kappa(x)$ is a local field of characteristic $p$ by [6, Lemma 3.17] and the corresponding homomorphism $C\to\kappa(x)$ gives a $G$-pseudocharacter $\Theta\in\mathrm{PC}_{G}^{\Gamma}(\kappa(x))$, such that $\varphi\circ\Theta\in\mathrm{PC}_{H}^{\Gamma}(\kappa(x))$ arises by scalar extension from the residual pseudocharacter attached to $\varphi\circ\overline{\rho}$.
For all $m\geq 0$, the map $(\varphi\circ\Theta)_{m}=\Theta_{m}\circ\varphi^{*}$ is the composition
$\mathcal{O}[H^{m}]^{H^{0}}\xrightarrow{\varphi^{*}}\mathcal{O}[G^{m}]^{G^{0}}\xrightarrow{\Theta_{m}}\mathrm{Map}(\Gamma^{m},\kappa(x))$
and has image in $\mathrm{Map}(\Gamma^{m},k)$. By 6.3 $\varphi^{*}$ is finite, so $\Theta_{m}$ has image in $\mathrm{Map}(\Gamma^{m},\overline{k})$. This implies that $\kappa(x)$ is algebraic over $k$ and contradicts the assumption that $x$ is a dimension $1$ point. ∎
@@
(37) $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$$G^{N}_{X^{\mathrm{ps}}_{G}}$$X^{\mathrm{gen}}_{H,\varphi\circ\overline{\rho}^{\mathrm{ss}}}$$H^{N}_{X^{\mathrm{ps}}_{H}}.$$\scriptstyle{f}$
@@ 8.3. Products
Recall that a $G$-valued pseudocharacter is a family of $\mathcal{O}$-algebra homomorphisms $\Theta_{m}:\mathcal{O}[G^{m}]^{G^{0}}\to\mathrm{Map}(\Gamma^{m},A)$ satisfying certain compatibility conditions. Restriction along the map $\mathcal{O}[G_{i}^{m}]^{G_{i}^{0}}\to\mathcal{O}[G^{m}]^{G^{0}}$ defines $G_{i}$-pseudocharacters $\Theta^{(i)}$ satisfying the same compatibility conditions; this can be directly seen from the description of pseudocharacters in terms of functors on the category of finitely generated free groups [50, Proposition 3.19]. Conversely given $\Theta^{(i)}\in\mathrm{PC}_{G_{i}}^{\Gamma}(A)$, we can define a $G$-pseudocharacter $\Theta^{(1)}\otimes\Theta^{(2)}$ by defining $(\Theta^{(1)}\otimes\Theta^{(2)})_{m}:\mathcal{O}[G^{m}]^{G^{0}}\to\mathrm{Map}(\Gamma^{m},A)$ as the tensor product $\Theta^{(1)}_{m}\otimes\Theta^{(2)}_{m}:\mathcal{O}[G_{1}^{m}]^{G_{1}^{0}}\otimes_{\mathcal{O}}\mathcal{O}[G_{2}^{m}]^{G_{2}^{0}}\to\mathrm{Map}(\Gamma^{m},A)$. This implies that the first map is bijective.
@@ 8.4. Functoriality in $\Gamma$
If $G=\operatorname{GL}_{d}$ then this is proved in [6, Lemma A.3]. In the general case, one could either use that $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\rightarrow R^{\mathrm{ps}}_{G}$ is finite, which follows from Propositions 6.4 and 7.4, or observe that the proof of [6, Lemma A.3] carries over to the more general setting. ∎
@@ 9.1. Trimming the component group
Since $G$ is generalised reductive $G/G^{0}$ is finite étale over $\mathcal{O}$ by Remark 2.6. The map $\mathfrak{S}\mapsto\mathfrak{S}(\overline{k})$ induces an equivalence of categories between finite étale $\mathcal{O}$-groups schemes and the category of finite groups together with $\operatorname{Gal}(\overline{k}/k)$-action by group automorphisms. Since $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}(k)$ is non-empty, $\Delta$ is contained in $(G/G^{0})(k)$ and hence the constant group scheme $\underline{\Delta}$ defined by $\Delta$ is a finite étale subgroup scheme of $G/G^{0}$.
@@
(39) $\nu:X^{\mathrm{gen},\tau}_{G}(A)\rightarrow(G/G^{0})_{X^{\mathrm{ps}}_{\operatorname{GL}_{d}}}^{N}(A),\quad\rho\mapsto(\pi_{G}(\rho(\gamma_{1})),\ldots,\pi_{G}(\rho(\gamma_{N}))),$
where $\gamma_{1},\ldots,\gamma_{N}$ is a generating set for $\Gamma$ chosen before 5.7. Since $\nu$ is $G^{0}$-equivariant it factors through $X^{\mathrm{git},\tau}_{G}\rightarrow(G/G^{0})^{N}_{X^{\mathrm{ps}}_{\operatorname{GL}_{d}}}$. We view an $N$-tuple $\underline{\delta}=(\delta_{1},\ldots,\delta_{N})\in\Delta^{N}$ as a closed constant subscheme of $\underline{\Delta}^{N}_{X^{\mathrm{ps}}_{\operatorname{GL}_{d}}}$ over $X^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ and define
@@
If $R$ is a complete local noetherian $\mathcal{O}$-algebra with residue field $k$ then the map $R\rightarrow k$ induces a bijection $(G/G^{0})(R)\overset{\sim}{\longrightarrow}(G/G^{0})(k)$ by [32, Exp. I, Cor. 6.2]. We apply this observation with $R=R^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ and let $\underline{\delta}\in(G/G^{0})^{N}(k)$ be the $N$-tuple corresponding to $\alpha:X^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}\rightarrow X^{\mathrm{git},\tau}_{G}\overset{\nu}{\rightarrow}(G/G^{0})^{N}_{X^{\mathrm{ps}}_{\operatorname{GL}_{d}}}$. Since the closed point in $X^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ maps to $(\pi_{G}(\overline{\rho}^{\mathrm{ss}}(\gamma_{1})),\ldots,\pi_{G}(\overline{\rho}^{\mathrm{ss}}(\gamma_{N})))\in(G/G^{0})^{N}_{X^{\mathrm{git},\tau}_{G}}(k)$ under $\alpha$ we conclude that $\underline{\delta}=(\pi_{G}(\overline{\rho}^{\mathrm{ss}}(\gamma_{1})),\ldots,\pi_{G}(\overline{\rho}^{\mathrm{ss}}(\gamma_{N})))$.
@@ 9.2. Absolutely irreducible locus
We say that a geometric point $x\in X^{\mathrm{gen}}_{G}$ is $G$-stable, if the orbit $G\cdot x$ is closed and $\dim Z_{G}(x)=\dim Z(G)_{\kappa(x)}$, where $Z_{G}(x)$ is the $G_{\kappa(x)}$-centraliser of $x$.
@@
The action of $G$ on $X^{\mathrm{gen}}_{G}$ factors through the action of $G/Z(G)$. The stability condition in the definition is equivalent to $G\cdot x$ closed and $\dim G\cdot x=\dim(G/Z(G))_{\kappa(x)}$. Hence, it is equivalent to $G/Z(G)$-stability in the sense of [57, Definition 1].
@@
Moreover, $H$ is reductive and $\dim Z_{G}(H)=\dim Z(G_{\kappa(x)})$.
@@
The equivalence of (1) and (2) follows from Corollary 6.11. The equivalence of (2) and (3) follows from Definition 2.26. The equivalence of (3) and (4) follows from [51, Proposition 16.7]. The last part follows from [42, Lemma 6.2]. ∎
@@
where the inequality in the middle follows from [42, Lemma 6.2], as the centre of $H_{\kappa(x)}$ is contained in the $L_{\kappa(x)}$-centraliser of $H_{\kappa(x)}$, and the other two follow from Lemma 2.2. ∎
@@
Let $\kappa$ be an algebraic closure of $\kappa(y)$. If $x\in X^{\mathrm{gen}}_{G,y}(\kappa)$ then the $G$-orbit $G\cdot x$ is closed, as $x$ is $G$-stable by Proposition 9.7. Proposition 6.7 implies that the $G^{0}$-orbit $G^{0}\cdot x$ is also closed. Since $X^{\mathrm{gen}}_{G,y}$ contains a unique closed $G^{0}$-orbit by Proposition 6.9 we conclude that $G^{0}(\kappa)$ acts transitively on $X^{\mathrm{gen}}_{G,y}(\kappa)$. Thus the orbit map
$G^{0}_{\kappa}/Z_{G^{0}_{\kappa}}(x)\rightarrow X^{\mathrm{gen}}_{G,y,\kappa}$
induces a bijection on $\kappa$-points. Since both schemes are of finite type over $\kappa$, we deduce that $\dim X^{\mathrm{gen}}_{G,y}=\dim G^{0}_{\kappa}/Z_{G^{0}_{\kappa}}(x)=\dim G^{0}_{\kappa}-\dim Z_{G^{0}_{\kappa}}(x)$. By passing to the underlying reduced subschemes of neutral components we obtain $(Z_{G^{0}_{\kappa}}(x)^{0})^{\mathrm{red}}=(Z_{G_{\kappa}}(x)^{0})^{\mathrm{red}}=(Z(G)_{\kappa}^{0})^{\mathrm{red}}$, where the last equality holds as $x$ is $G$-stable. Thus $\dim Z_{G^{0}_{\kappa}}(x)=\dim Z(G)_{\kappa}$ and we conclude that
@@
Corollary 9.9 says that $\dim V_{y}=d$, where $d:=\dim G_{k}-\dim Z(G_{k})$, for all $y\in U$. The inequality $\dim V\leq\dim U+d$ follows directly from [6, Lemma 3.18 (6)] (or alternatively [59, Tag 0BAG]).
To prove the reverse inequality we may assume that $U$ is irreducible and, since $V\rightarrow U$ is surjective by [57, Theorem 3 (ii)], we may replace $V$ by an irreducible component, whose image contains the generic point of $U$. Lemma 2.1 in [6] implies that $V$ is $G^{0}$-invariant and [57, Theorem 3 (iii)] implies that the image of $V$ in $U$ is closed, and hence $V\rightarrow U$ is surjective. It follows from [59, Tag 02JU] applied with $s\in U$ a closed point such that $\dim U=\dim\mathcal{O}_{U,s}$ and $x\in V$ a closed point, which maps to $s$, that $\dim V\geq\dim U+d$. ∎
@@
Remark 9.15.
The referee suggested getting around the problem explained in 9.13 and 9.14 by working only with $G$-normalisers of parabolic subgroups of $G^{0}$ (which are $R$-parabolic by [3, Proposition 6.1]) instead of all $R$-parabolic subgroups. This definition would save some work in Section 9.1, but would not match the standard definition for $G$-irreducibility in the literature such as [3], which we use. If $\pi_{G}\circ\overline{\rho}^{\mathrm{ss}}$ is surjective then both definitions of irreducibility coincide.
@@ 10. Bounds for the dimensions of reductive subgroups
Let $G$ be a generalised reductive group over an an algebraically closed field $\kappa$. If $H$ is a closed reduced subgroup scheme of $G$, such that the unipotent radical of $H^{0}$ is trivial then $H$ is also generalised reductive by 2.7. Let $G^{\prime}$ denote the derived subgroup of $G^{0}$. Then $G^{\prime}$ is semisimple and $G^{0}=Z(G^{0})G^{\prime}$, [10, 14.2]. Moreover, $G^{\prime}$ is equal to its own derived subgroup, [17, Example 1.1.16].
@@
After replacing $G$ by $G^{\prime}$ we may assume that $G$ is connected and semisimple. One may check that the unipotent radical of $(G^{\prime}\cap H)^{0}$ is normal in $H^{0}$. Since $H^{0}$ is reductive, $(G^{\prime}\cap H)^{0}$ is also reductive. Thus after replacing $H$ by $H\cap G^{\prime}$ we may assume that $H$ is closed reductive subgroup of $G$.
Let $B_{0}=T_{0}U_{0}$ be a Borel subgroup of $H$ with the unipotent radical $U_{0}$ and a maximal torus $T_{0}$. Let $B=TU$ be a Borel subgroup of $G$ with the unipotent radical $U$ and a maximal torus $T$, such that $U$ contains $U_{0}$ and $T$ contains $T_{0}$. We have $\dim G=2\dim U+\dim T$ and $\dim H=2\dim U_{0}+\dim T_{0}$. If $U\neq U_{0}$ then $\dim U>\dim U_{0}$ and we deduce that $\dim G-\dim H\geq 2$.
Let us assume that $U=U_{0}$ and let $U^{-}_{0}$ be the unipotent radical of a Borel subgroup of $H$, which is opposite to $B_{0}$ with respect to $T_{0}$. Since $U_{0}=U$, $U_{0}^{-}$ is a unipotent radical of a Borel subgroup of $G$. Since $U_{0}^{-}\cap U_{0}=\{e\}$, [9, Proposition 4.10 e)] implies that $U^{-}_{0}$ is a unipotent radical of a parabolic subgroup of $G$, which is opposite to $B$ with respect to $T$. Since $G$ is by assumption semisimple, $U_{0}$ and $U_{0}^{-}$ generate it as a group by [58, Theorem 8.1.5(i)]. Thus $G=H$. ∎
@@
If $\dim T=1$ then $G^{\prime}=\operatorname{SL}_{2}$ by [58, 7.2.4], as the assumption that $G^{\prime}$ is simply connected rules out $\operatorname{PGL}_{2}$. If $\mathrm{char}(\kappa)\neq 2$ then the action of $G^{\prime}$ on $\operatorname{Lie}G^{\prime}$ is an irreducible representation, isomorphic to $\operatorname{Sym}^{2}(\mathrm{Std})$, where $\mathrm{Std}$ is the standard $2$-dimensional representation of $\operatorname{SL}_{2}$. If $\mathrm{char}(\kappa)=2$ then $\operatorname{Lie}G^{\prime}$ is a non-split extension $0\rightarrow\mathbf{1}\rightarrow\operatorname{Lie}G^{\prime}\rightarrow V\rightarrow 0$, with $\mathbf{1}$ equal to the subspace of scalar matrices and $V$ an irreducible $2$-dimensional representation. In both cases we obtain the claim.
Let $\dim T$ be arbitrary and let $\chi:\operatorname{Lie}G^{\prime}\rightarrow\mathbf{1}$ be $G^{\prime}$-invariant. Let us abbreviate $\mathfrak{g}=\operatorname{Lie}G^{\prime}$. We have $\mathfrak{g}=\mathfrak{g}^{T}\oplus\bigoplus\nolimits_{\alpha\in\Phi}\mathfrak{g}_{\alpha}$, where $\Phi$ is the set of roots, and $\mathfrak{g}^{T}=\operatorname{Lie}(T)$, [10, 13.18(1)]. Let $\alpha\in\Phi$ and let $G_{\alpha}$ be the centraliser of $(\operatorname{Ker}\alpha)^{0}$ in $G^{\prime}$. Then $G_{\alpha}$ is reductive of semisimple rank $1$, [10, 13.18(4)]. Let $G^{\prime}_{\alpha}$ be the derived subgroup of $G_{\alpha}$. The assumption that $G^{\prime}$ is simply connected implies that $G^{\prime}_{\alpha}\cong\operatorname{SL}_{2}$ by [17, 5.3.9]. It follows from [17, 1.2.7] that $\operatorname{Lie}G_{\alpha}^{\prime}=\mathfrak{g}_{-\alpha}\oplus\operatorname{Lie}(\alpha^{\vee}(\mathbb{G}_{m}))\oplus\mathfrak{g}_{\alpha}$. The restriction of $\chi$ to $\operatorname{Lie}G^{\prime}_{\alpha}$ is $G^{\prime}_{\alpha}$-invariant and hence zero by the previous part. Hence, $\chi$ is zero on $\operatorname{Lie}(\alpha^{\vee}(\mathbb{G}_{m}))$ and $\mathfrak{g}_{\alpha}$ for every coroot $\alpha^{\vee}$.
@@
Let $P$ be an $R$-parabolic subgroup of $G$, let $U$ its unipotent radical and let $L$ be an $R$-Levi subgroup of $P$. Then $P^{0}=P\cap G^{0}$ is a parabolic subgroup of $G^{0}$ and $L^{0}=L\cap G^{0}$ is a Levi subgroup of $P^{0}$, [42, Proposition 5.2]. In particular, $L^{0}$ is reductive. We let $L^{\prime}$ be the derived subgroup of $L^{0}$.